OpenStudy (anonymous):

I worked out this problem however, my answer is wrong. What did I do wrong? 5x^(2/3) (x^(1/3) + x^(-2/3)) ³√((5x)²) × (x^(1/3) + x^(-2/3)) ³√(25x³) + (³√(25x²))/(³√(x²)) My answer: x ³√(25) + ³√(25) Correct Answer: 5x + 5

OpenStudy (akashdeepdeb):

\[5x^{2/3}(x^{1/3} + x^{-2/3})\] Open the parenthesis. You'll get: \[5x^{2/3}*x^{1/3} + 5x^{2/3}*x^{-2/3})\] Which further gives us. \[5x^{\frac{2}{3} + \frac{1}{3}} + 5x^{\frac{2}{3} - \frac{2}{3}}\] Using the simple exponent rules. We have: \[5x^1 + 5x^0 = 5x + 5 \] Understood? :)

OpenStudy (triciaal):

is there a parenthesis after the first 5?

OpenStudy (jdoe0001):

\(\Large \bf 5x^{\frac{2}{3}}x^{\frac{1}{3}}+x^{-\frac{2}{3}}\qquad or\qquad (5x)^{\frac{2}{3}}x^{\frac{1}{3}}+x^{-\frac{2}{3}}\quad ?\)

OpenStudy (anonymous):

@triciaal No, there isn't. @AkashdeepDeb That makes much more sense, thank you very much! :D @jdoe0001 It's more like the second one. Thank you everyone for helping! :D

OpenStudy (akashdeepdeb):

:)